OpenStudy (anonymous):

Anyone that answers this i will be a fan 15x-8-3x=16

OpenStudy (anonymous):

a ceiling fan?

jigglypuff314 (jigglypuff314):

15x-3x = (15-3)x = 12x so 15x-8-3x = 16 -> 12x - 8 = 16 then add 8 to both sides +8 +8 12x = 24 then divide both sides by 12 /12 /12 so x = 24 divided by 12 = ?

OpenStudy (anonymous):

What is 24/12?